OpenStudy (anonymous):

What is the graph of the function f(x) = the quantity of negative x squared minus 2 x minus 2, all over x minus 2? **MEDAL/FAN**

OpenStudy (anonymous):

@blackeistien238821 @rishavraj @Legends

OpenStudy (anonymous):



OpenStudy (anonymous):

@wil476003 @Quanta________ @jim_thompson5910

jimthompson5910 (jim_thompson5910):

There are two ways you can do this 1) plot a bunch of points and connect them with a curve 2) use a graphing calculator like TI83 or something like desmos https://www.desmos.com/calculator

jimthompson5910 (jim_thompson5910):

if you use method 1), then you plug in various values of x to get the corresponding values of y example: if x = 0, then \[\Large y = \frac{-x^2-2x-2}{x-2}\] \[\Large y = \frac{-0^2-2(0)-2}{0-2}\] \[\Large y = 1\] so the point (0,1) lies on the function curve. To find other points, you follow the same steps for different x values. The only thing not allowed is x = 2.

OpenStudy (anonymous):

Ahh! Okay! Thank-you<3

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

How would the graph look on desmos because it isn't workinf for me